OpenStudy (anonymous):

please help integral 1 to 2 ((3)/(sqrt(x)sqrt(4-x)))dx...

OpenStudy (anonymous):

\[ \int\limits_{2}^{1}(3)/(\sqrt(x)\sqrt(4-x)\] means this????????

OpenStudy (anonymous):

2 on top 1 on bottom

OpenStudy (anonymous):

@paitucker

OpenStudy (anonymous):